A standard 1 kilogram weight is a cylinder 50.5 mm in height and 46.5 mm in diameter. What is the density of the material? kg/m3
Answer #1

Density = mass/ volume

Now volume of cylinder is v = (22/7)×r^2×h

= 22/7×23.25×23.25×50.5

= 85,795 mm^3

Now volume in cubic meter V =8.5795 ×10^(-5) m^3

So density d = m/V = 1/(8.5795×10^(-5))

Or d = 1.166× 10^4 kg/m^3